# A Comprehensive Survey of Artificial Intelligence Techniques for Talent Analytics

### v2

@@ 2.1.3 Organizational Data
Job Search Websites  [246, 245, 31, 357, 318]
2.2 External Data
Social Media: Widely-used social media platforms that contribute to a comprehensive understanding of the labor market include Twitter  [^1] https://twitter.com, Facebook  [^2] https://www.facebook.com, and news reports. With the help of NLP and Topic Model techniques [49], numerous studies have been carried out to explore the semantic information in this corpus. For example, more than 60,000 tweets related to nine energy companies were collected for sentiment analysis expressed on Twitter [188]. To gain further insight into the impact of public opinion, Spears et al. [376] collected earnings reports and news articles spanning eight years from four companies. The results indicate that companies may face a decline in valuation when they receive negative publicity.
Job Search Websites: Recent years have witnessed the rapid growth of job search websites, such as Indeed  [^3] https://www.indeed.com, LinkedIn  [^4] https://www.linkedin.com, and Glassdoor  [^5] https://www.glassdoor.com. Specifically, Indeed and Glassdoor allow users to comment on a company, providing an overall understanding of the employer brand. For instance, Lin et al. [246, 245] collaboratively modeled both textual (i.e., reviews) and numerical information (i.e., salaries and ratings) for learning latent structural patterns of employer brands. In addition, Bajpai [31] leveraged the data from Glassdoor to perform aspect-level sentiment analysis. Along this line, large-scale reviews of Fortune 500 companies are collected to identify topics that matter to employees  [357]. Differently, LinkedIn provides a wide range of business services, including job listings, professional profile creation, and career development services, with personal profiles being the most analyzed, as they describe users’ employment history. For example, Park et al. [318] used LinkedIn’s employment history data from more than 500 million users over 25 years to construct a labor flow network of over 4 million firms worldwide, demonstrating a strong association between the influx of educated workers and financial performance in detected geo-industrial clusters.
Furthermore, there are also several third-party business investigation platforms that offer detailed information about companies and their board members’ relationships, such as Crunchbase  [^6] https://www.crunchbase.com, Owler  [^7] https://www.owler.com, Tianyancha  [^8] https://www.tianyancha.com and Aiqicha  [^9] https://www.aiqicha.com. These details can be viewed as complementary information to the job search websites. Building upon this foundation, one can gain deeper insights into the aligned companies and conduct more relevant research, such as analyzing cooperative and competitive relationships [96] and providing investment target recommendations [81].
2.3 Data Processing
2.3.1 Data Collection
Internel Data. The Current business environment is typically dependent on data systems [219]. Internal data are collected from the internal enterprise management systems, which are known as enterprise resource planning (ERP) system [362], customer relationship management (CRM) system [127] and applicant tracking system (ATS) system [300]. ERP systems are comprehensive business management tools integrating various functions such as finance, sales, materials management, HR, production planning, and supply chain. CRM systems facilitate customer interaction and communication, encompassing customer information management, sales opportunities, and customer service. An ATS is computer software that human resource departments use to process the overwhelming number of applications they receive for job openings. These systems orderly store recruitment, employee, and organizational data, facilitating efficient collection and processing.
Externel Data. As online services rapidly evolve, a growing number of individuals are turning to social media and job search websites to exchange job-related information and explore employment opportunities. These interactive platforms host a vast array of talent information due to their extensive user base. Additionally, third-party business investigation platforms provide detailed insights into companies and the relationships of their board members. Data acquisition methods on these platforms vary. Third-party data collection websites typically aggregate data from their participating members. Meanwhile, web crawlers [215] can extract rich information from website pages, provided legal and regulatory compliance is ensured. Moreover, many job search websites retain substantial amounts of user data that isn’t publicly disclosed. Typically, this data can be utilized for scientific research purposes following encryption and other privacy safeguards.
2.3.2 Data Preprocessing
Structured Data. Structured data, in simple terms, is a database such as ERP system [120], that has a standardized format for efficient access by software and humans alike. Besides the information of employees and companies are orderly collected into the database. On some data collection websites, a lot of user information is also strictly stored through databases, such as records of interactions between users and the platform, including clicking and browsing, etc. Since these data have been stored in a standardized format, filtering and integration of corresponding data can generally be achieved by connecting different tables and setting data filtering conditions on the tables [32].
Unstructured Data. Unstructured data [396] refers to information that does not conform to the conventional row-column structure found in traditional databases. Within the labor market context, data primarily manifests as text, comprising a blend of structured and unstructured fields. Structured fields denote specific categories like job titles (e.g., occupation), location, etc., while unstructured fields provide a broader description of vacancy content. Approximately 80% of data held by firms today is unstructured [35], expanding at a rate fifteen times faster than structured data. While some approaches skip the processing of natural text, opting instead for direct classification tasks on such text, as observed in the classification of web job vacancies [55], others require extraction and subsequent processing of information from unstructured data into structured data. For instance, extracting skill information from job descriptions [83] involves the identification of skill words through regular expression matching. These identified words are then subjected to expert evaluation for further refinement, resulting in a set of relevant skill words for each job description. Subsequent statistical analysis across all jobs yields a comprehensive frequency distribution of skill words as the structured data, facilitating downstream tasks such as skill demand forecasting.
2.3.3 Data Cleaning and Debias
Data Quality Issues. There are various data quality issues: missing data, duplicated data, extraneous data, and inconsistent data [93]. These issues will introduce biases into analyses and lead to inaccurate conclusions. We introduced these issues as follows:
TABLE II: The table of collected talent analytic-related open datasets.
@@ 2.3.3 Data Cleaning and Debias
Cleaning and Debias Methods. Data cleaning and debias is an iterative process tailored to the requirements and semantics of specific analysis tasks [218]. This process consists of transforming raw data into consistent data that can be analyzed. Herein, we delineate several frequently employed methodologies in talent analytics for data cleaning and debias.
2.3.4 Limitations
  • Lack of benchmark datasets: Most current research in talent analytics heavily depends on proprietary data related to job applicants, employees, and organizations. The sensitive nature of this data often prevents it from being publicly accessible, severely restricting the development of publicly available benchmark datasets. This limitation impedes the standardization of problem definitions and comparative analyses of methodologies, thus slowing down technological advancements in the field. While some job-related data has been made open-source, it remains limited in scale and temporal coverage. Consequently, there is an urgent need to create comprehensive open-source benchmark datasets for pivotal tasks within this field. We present several open datasets, as shown in Table II, primarily related to resume parsing and employee turnover prediction. These datasets can help researchers in developing standardized datasets for uniform comparisons. Additionally, the data anonymization and de-identification methods require more extensive consideration to facilitate the construction of more open-source datasets for various scenarios.
  • Diversity deficiencies in datasets: The datasets commonly used in AI-based talent analytics often lack adequate consideration of diverse populations, especially underrepresented minorities. This deficiency hampers efforts to evaluate and ensure algorithmic fairness. Models developed from such datasets may exhibit biases, leading to inequitable outcomes among various demographic groups. Therefore, it is essential to integrate more diverse datasets to promote fairness and enhance the overall performance of AI algorithms in talent analytics.
3 Talent Management
Talent management, which focuses on placing the right person in the right job at the right time, has emerged as a predominant human capital topic in the early twenty-first century [66, 63]. Specifically, the management process includes the whole process of talent entering the organization to development. Accordingly, first, we describe various intelligent talent recruitment scenarios, such as job posting generation [254, 332] and talent searching [162, 274, 160]. After entering the organization, in order to ensure the sustainable development of talent, timely and accurate feedback is critical. As a result, second, we discuss two primary issues in talent assessment: interview question recommendation [363, 336, 365] and assessment scoring [304, 171, 77, 234]. Finally, from the whole development process, career development is important for managing human capital resources and individual development. Accordingly, we outline several post-employment career development problems, including course recommendations [377, 402, 401] for employee training and employee dynamics analysis [477, 259, 386]. In the following sections, we will delve into these issues in detail.
TABLE III: The table of collected papers related to talent management.

### v3

@@ B.2 Data Preprocessing
Job Search Websites  [27, 28, 68, 69, 70][71]
TABLE S2: The table of collected talent analytics-related open datasets.
@@
Structured Data. Structured data, in simple terms, is a database such as ERP system [313], that has a standardized format for efficient access by software and humans alike. Besides, the information of employees and companies is orderly collected into the database. On some data collection websites, a lot of user information is also strictly stored through databases, such as records of interactions between users and the platform, including clicking and browsing. Since these data have been stored in a standardized format, filtering and integration of corresponding data can generally be achieved by connecting different tables and setting data filtering conditions on the tables [314].
Unstructured Data. Unstructured data [321] refers to information that does not conform to the conventional row-column structure found in traditional databases. Within the labor market context, data primarily manifests as text, comprising a blend of structured and unstructured fields. Structured fields denote specific categories like job titles (e.g., occupation), location, etc., while unstructured fields provide a broader description of vacancy content. Approximately 80% of data held by firms today is unstructured [322], expanding at a rate fifteen times faster than structured data. While some approaches skip the processing of natural text, opting instead for direct classification tasks on such text, as observed in the classification of web job vacancies [323], others require extraction and subsequent processing of information from unstructured data into structured data. For instance, extracting skill information from job descriptions [86] involves the identification of skill words through regular expression matching. These identified words are then subjected to expert evaluation for further refinement, resulting in a set of relevant skill words for each job description. Subsequent statistical analysis across all jobs yields a comprehensive frequency distribution of skill words as structured data, facilitating downstream tasks such as skill demand forecasting.
B.3 Data Cleaning and Debias
Data Quality Issues. There are various data quality issues: missing data, duplicated data, extraneous data, and inconsistent data [324]. These issues will introduce biases into analyses and lead to inaccurate conclusions. We introduce these issues as follows:
TABLE S3: The table of collected papers related to talent management.